OpenStudy (anonymous):

What is the value of X/Y If x=9/4 and y=3/5 ?

jhonyy9 (jhonyy9):

9/4/3/5 = 9/4 *5/3 = 45/12 = 15/4

OpenStudy (anonymous):

Thanks

jhonyy9 (jhonyy9):

yw good luck bye